OpenStudy (anonymous):

find the slope intercept form for the line satisfying the following conditions: x intercept 3, y intercept 2 - 5

OpenStudy (chihiroasleaf):

the line through points \[\left( 0,\frac{ 2 }{ 5 } \right)\] and \[\left( 3,0 \right)\] so, the slope of the lines is \[\frac{\frac{ 2 }{ 5 } - 0 }{ 0 -3 }=-\frac{ 2 }{ 15 }\] the y-intercept is \[\frac{ 2 }{ 5 }\] so, the slope-intercept form of the equation line is \[y=-\frac{ 2 }{ 15 }x+\frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

Thank you for the help and explanation. :)

OpenStudy (chihiroasleaf):

you're welcome... :)